# Differentiable Variable Fonts

### v1

@@ 1 Introduction
To address this disconnect between graphics tools and typography, we build on a new font technology: variable fonts. A variable font is a single font file that stores a continuous range of design variants. Instead of offering a fixed style, a variable font exposes one or more parameters, called axes, that control stylistic variation. Each axis corresponds to a numerical value that can be adjusted independently and may correspond to conventional typographic properties such as weight or width, or to custom dimensions defined by the type designer (See Fig. 1). This continuous design space is powerful – as long as the text manipulation is constrained to a variable font, it will be legible and consistent in style. However, variable fonts are underutilized in creative applications because artists need to manually adjust each parameter in this often high-dimensional space.
We formalize the variable font mechanism in a compact mathematical model, revealing the highly non-linear, non-invertible mapping of variable font parameters to the underlying vector graphics representing the text (Section 3). Based on our formulation, we implement variable font interpolation in a way that exposes gradients with respect to a font’s axes, enabling gradient-based optimization directly in the variable font space (Section 4). This makes it possible to cast a broad range of graphics editing tasks as energy minimization problems defined directly over variable font parameters.
@@ 2.1 Parametric Font Representations
Alongside industrial developments, academic research has explored parametric representations of fonts—ranging from early work that treated high-level glyph features like stems and serifs [SR98] or reusable components [HH01] as parameters, to more recent approaches that learn continuous font spaces from examples to support interpolation [CK14]. While these methods can be effective for font generation, they often offer limited user control over the design space — an important requirement in applications like logo design, title sequences, and motion graphics, where maintaining stylistic consistency is critical for brand identity. Variable fonts provide a practical alternative: they define a constrained, interpretable space crafted by type designers, allowing controlled manipulation without sacrificing design intent. They are also supported across all major browsers and enable efficient optimization, making them well-suited for deployment in real-world design tools.
@@ 2.2 Font Manipulation
Digital fonts represent each glyph’s geometry as a collection of Bézier curves. In standard design workflows, text is converted to vector paths and edited using general-purpose tools such as Adobe Illustrator, Figma, Inkscape, or Affinity Designer. This detaches the text from its typographic structure, reducing it to unstructured vector geometry. While this provides flexibility, making precise edits requires multiple iterations of careful curve manipulation to preserve legibility and aesthetic appeal. The process is time-consuming and demands expert judgment. Likewise, automatic shape manipulation techniques from geometry processing operate on vector graphics or meshes without regard for typographic structure, and can similarly compromise legibility when applied to text. Moreover, many such techniques aim to preserve area, length, or angles under deformation [IMH05, SA07, WXW06, CPSS10, WG10, SBCBG11], making them ill-suited for common typographic modifications, such as increasing stroke weight or applying slant, which alter area and internal angles in ways these methods are not designed to accommodate. (See Fig. 2)
Attempts to preserve meaningful structure during 2D or 3D shape editing have a long history across various application domains. Some methods rely on user-defined constraints [Gle92, HLW93, Sut64], while others attempt to automatically detect and preserve salient structures [CLDD09, BL15, AVR22]. In contrast, variable fonts are carefully constructed by type designers to ensure that every point in the design space yields a legible and stylistically coherent instance. By allowing the end user to select an appropriate variable font, we can take advantage of this expert-authored structure, with the assurance that edits confined to the font’s defined variation space will preserve legibility.
@@ 2.3 Kinetic Typography
More recently, image-based approaches have been explored [LMO24, PBSJ24, MLTX19]. Wakey-Wakey [XZY23] performs example-based motion transfer by optimizing control points of text vector outlines. However, such methods either neglect core principles of typography—legibility and stylistic consistency—or rely on heuristic regularization, which still cannot guarantee their preservation. In contrast, our work allows animating text directly in the space of variable fonts, where legibility and stylistic coherence are preserved by construction. To drive motion, we employ variational methods for physics based animation.
@@ 2.4 Font Matching
Font recognition is the task of identifying a typeface given an image of a text fragment. It has been studied extensively, with approaches ranging from nearest-class-mean classifiers with learned features [CYJ14] to deep learning–based methods [WYJ15a, WYJ15b, WLTX18]. These methods aim to select the closest typeface from a large library of static fonts. In contrast, our objective is to recover the best-matching instance within a variable font, operating in a continuous parameter space rather than over discrete classes. To this end, we combine differentiable rasterization [LLMRK20] with our differentiable variable font framework to directly optimize font variation axes against a simple image loss—without requiring any learning.
@@ 3 Variable Font Mathematics
Variable Fonts are a generalization of earlier efforts like Adobe’s Multiple Master and Apple’s GX fonts. Their ability to easily interpolate between designer-specified glyphs belies complex underlying interpolation modes expressed through a combination of pseudocode and descriptive text, distributed across several sections of the OpenType documentation. In this section we distill this implementation-skewed schema to a compact mathematical formulation of variable font interpolation. This mathematical foundation enables the development of a differentiable variable font framework, which can then be applied to support new modes of interaction and control.
@@
Each variable axis $i$ is defined over a numeric domain with a minimum, maximum, and default value: $s_{i}^{min}\leq s_{i}\leq s_{i}^{max}$. An instance of the variable font is specified by a vector $\mathbf{s}=(s_{1},s_{2},\ldots,s_{n})\in\mathbb{R}^{n}$, where $s_{i}$ is chosen (by the user— via axis sliders in a UI) within the axis’s allowed range. Internally, these axis values are normalized to a standard coordinate system via a piecewise-linear transformation, which maps $\{s^{\text{min}}_{i},s^{0}_{i},s^{\text{max}}_{i}\}$ to $\{-1,0,1\}$.
@@
where $\Theta=[\mathbf{w}_{1}\mathbf{w}_{2}\ldots\mathbf{w}_{l}]$ are the parameters for each glyph in the word and $\overleftarrow{\mathbf{p}}^{\text{ }g}$, $\overrightarrow{\mathbf{p}}^{\text{ }g}$ denotes the left and right sidebearing points of $g$ respectively. Note that in this formula, we abuse notation slightly by assuming integer glyph position will index the correct glyph in the font.
@@ 4 Optimization of Variable Fonts
Axis limit constraints. The axis weights $\Theta\in[-1,1]^{n\cdot l}$ must remain within their valid bounds, since interpolation is undefined outside these limits. In practice, we clamp any out-of-range values before evaluating the interpolation to ensure it is well-defined. However, clamping alone creates regions with zero gradient. To address this, we project $\Theta$ back within the valid bounds after each update.
@@ 5 Applications
We now demonstrate the utility of our differentiable variable font optimization framework through several applications. These include interactive editing tools (Section 5.1, 5.2), physics driven kinetic typography (Section 5.3), and automatic font matching (Section 5.4). Each of these applications builds on the same underlying formulation but defines different energy objectives tailored to the task.
@@ 5.1 Direct Manipulation
where the second term regularizes the solution to remain close to the previous font instance. Additional editing constraints, such as fixing points, enforcing shared $x$ or $y$ coordinates, or encouraging or maintaining collinearity, can be incorporated naturally as extra least-squares terms in the same formulation (See Fig. 9). We solve this optimization using the Levenberg–Marquardt algorithm.
@@ 5.2 Overlap Aware Modeling
Our differentiable framework automates this process by updating the axis weights to resolve overlaps efficiently. (See Fig. 10). Collision detection is performed by densely sampling points along all glyph outlines and any background objects. For each colliding point, we find the closest point on the other surface and its normal, defining a unilateral collision plane. Let $\mathbf{t}$ contain the curve indices and Bézier parameters for all points that need to be projected, and let $\mathbf{p}(\Theta,\mathbf{t})$ give the corresponding stacked positions under the current axis weights. Similarly, let $\mathbf{b}$ and $\mathbf{N}$ be the stacked closest points and normals. We define the collision energy as a differentiable least-squares objective:
@@ 5.3 Physics Driven Kinetic Typography
For time integration, we adopt an iterative constraint projection approach inspired by projective dynamics [BML14]. Unlike standard projective dynamics, however, we cannot rely on a global linear solve since the collision constraint projection is nonlinear. Instead, we iteratively minimize the combined elastic, kinetic, and collision energies using the Levenberg–Marquardt optimizer at each timestep.
@@ 5.4 Font Matching
using gradient descent. In practice, we use differentiable vector graphics rasterization [LLMRK20] and employ an Adam optimizer (with 100 iterations), which reliably converges to visually accurate matches without additional tuning (Fig. 12).

### v2

@@ 1 Introduction
To address this disconnect between graphics tools and typography, we build on a recent font technology: variable fonts. A variable font is a single font file that stores a continuous range of design variants. Instead of offering a fixed style, a variable font exposes one or more parameters, called axes, that control stylistic variation. Each axis corresponds to a numerical value that can be adjusted independently and may correspond to conventional typographic properties such as weight or width, or to custom dimensions defined by the type designer (See Fig. 1). This continuous design space is powerful – as long as the text manipulation is constrained to a variable font, it will be legible and consistent in style. However, variable fonts are underutilized in creative applications because artists need to manually adjust each parameter in this often high-dimensional space.
We formalize the existing OpenType variable font interpolation mechanism in a compact mathematical model, revealing the highly non-linear, non-invertible mapping of variable font parameters to the underlying vector graphics representing the text (Section 3). Based on our formulation, we implement variable font interpolation in a way that exposes gradients with respect to a font’s axes, enabling gradient-based optimization directly in the variable font space (Section 4). This makes it possible to cast a broad range of graphics editing tasks as energy minimization problems defined directly over variable font parameters.
@@ 2.1 Parametric Font Representations
Alongside industrial developments, academic research has explored parametric representations of fonts—ranging from early work that treated high-level glyph features like stems and serifs [SR98] or reusable components [HH01] as parameters, to more recent approaches that learn continuous font spaces from examples to support interpolation [CK14]. While these methods can be effective for font generation, they often offer limited user control over the design space — an important requirement in applications like logo design, title sequences, and motion graphics, where maintaining stylistic consistency is critical for brand identity. Variable fonts provide a practical alternative: they define a constrained, interpretable space crafted by type designers, allowing controlled manipulation without sacrificing design intent. They are also supported across all major browsers and enable efficient optimization, making them well-suited for deployment in real-world design tools. Rather than proposing a new font representation, our contribution lies in providing a concise mathematical formalization of the existing OpenType variable font mechanism and making it differentiable, enabling gradient-based optimization.
@@ 2.2 Font Manipulation
Digital fonts represent each glyph’s geometry as a collection of Bézier curves. In standard design workflows, text is converted to vector paths and edited using general-purpose tools such as Adobe Illustrator, Figma, Inkscape, or Affinity Designer. This detaches the text from its typographic structure, reducing it to unstructured vector geometry. While this provides flexibility, making precise edits requires multiple iterations of careful curve manipulation to preserve legibility and aesthetic appeal. The process is time-consuming and demands expert judgment. Likewise, automatic shape manipulation techniques from geometry processing operate on vector graphics or meshes without regard for typographic structure, and can similarly compromise legibility when applied to text. Moreover, many such techniques aim to preserve area, length, or angles under deformation [IMH05, SA07, WXW*06, CPSS10, WG10, SBBG11], making them ill-suited for common typographic modifications, such as increasing stroke weight or applying slant, which alter area and internal angles in ways these methods are not designed to accommodate. (See Fig. 2)
Attempts to preserve meaningful structure during 2D or 3D shape editing have a long history across various application domains. Some methods rely on user-defined constraints [Gle92, HLW93, Sut64], while others attempt to automatically detect and preserve salient structures [CLDD09, BL15, AVR*22]. In contrast, variable fonts are carefully constructed by type designers to ensure that every point in the design space yields a legible and stylistically coherent instance. By allowing the end user to select an appropriate variable font, we can take advantage of this expert-authored structure, with the assurance that edits confined to the font’s defined variation space will preserve legibility.
@@ 2.3 Kinetic Typography
More recently, image-based approaches have been explored [LMO*24, PBSJ24, MLTX19]. Wakey-Wakey [XZY*23] performs example-based motion transfer by optimizing control points of text vector outlines. However, such methods either neglect core principles of typography, legibility and stylistic consistency, or rely on heuristic regularization, which still cannot guarantee their preservation. In contrast, our work allows animating text directly in the space of variable fonts, where legibility and stylistic coherence are preserved by construction. To drive motion, we employ variational methods for physics based animation.
@@ 2.4 Font Matching
Font recognition is the task of identifying a typeface given an image of a text fragment. It has been studied extensively, with approaches ranging from nearest-class-mean classifiers with learned features [CYJ*14] to deep learning–based methods [WYJ*15, WYJ*15a, WLTX18]. These methods aim to select the closest typeface from a large library of static fonts. In contrast, our objective is to recover the best-matching instance within a variable font, operating in a continuous parameter space rather than over discrete classes. To this end, we combine differentiable rasterization [LLMR20] with our differentiable variable font framework to directly optimize font variation axes against a simple image loss without requiring any learning.
@@ 3 Variable Font Mathematics
Variable Fonts, as defined by the OpenType specification, are a generalization of earlier efforts like Adobe’s Multiple Master and Apple’s GX fonts. Their ability to easily interpolate between designer-specified glyphs belies complex underlying interpolation modes expressed through a combination of pseudocode and descriptive text, distributed across several sections of the OpenType documentation. In this section we distill this implementation-skewed schema to a compact mathematical formulation of variable font interpolation. This mathematical foundation enables the development of a differentiable variable font framework, which can then be applied to support new modes of interaction and control.
@@
Each variable axis $i$ is defined over a numeric domain with a minimum, maximum, and default value: $s_{i}^{min}\leq s_{i}\leq s_{i}^{max}$. An instance of the variable font is specified by a vector $\mathbf{s}=(s_{1},s_{2},\ldots,s_{n})\in\mathbb{R}^{n}$, where $s_{i}$ is chosen (by the user via axis sliders in a UI) within the axis’s allowed range. Internally, these axis values are normalized to a standard coordinate system via a piecewise-linear transformation, which maps $\{s^{\text{min}}_{i},s^{0}_{i},s^{\text{max}}_{i}\}$ to $\{-1,0,1\}$.
@@
where $\Theta=[\mathbf{w}^{1}\mathbf{w}^{2}\ldots\mathbf{w}^{l}]$ are the parameters for each glyph in the word and $\overleftarrow{\mathbf{p}}^{\text{ }g}$, $\overrightarrow{\mathbf{p}}^{\text{ }g}$ denotes the left and right sidebearing points of $g$ respectively. Note that in this formula, we abuse notation slightly by assuming integer glyph position will index the correct glyph in the font.
@@ 4 Optimization of Variable Fonts
Axis limit constraints. The axis weights $\Theta\in[-1,1]^{n\cdot l}$ must remain within their valid bounds, since interpolation is undefined outside these limits. In practice, we clamp any out-of-range values before evaluating the interpolation to ensure it is well-defined. However, clamping alone creates regions with zero gradient. To address this, we project $\Theta$ back to the valid bounds after each update.
@@ 5 Applications
We now demonstrate the utility of our differentiable variable font optimization framework through several applications. These include interactive editing tools (Section 5.1, 5.2), physics driven kinetic typography (Section 5.3), and automatic font matching (Section 5.4). Each of these applications builds on the same underlying formulation but defines different energy objectives tailored to the task. We also report representative runtime statistics for these applications in Section 5.5.
@@ 5.1 Direct Manipulation
where the second term regularizes the solution to remain close to the previous font instance. Additional editing constraints, such as fixing points, enforcing shared $x$ or $y$ coordinates, or encouraging or maintaining collinearity, can be incorporated naturally as extra least-squares terms in the same formulation (See Fig. 9, video 2:35–2:57). We solve this optimization using the Levenberg–Marquardt algorithm. We conducted a pilot user study to obtain preliminary feedback on the usability of direct manipulation compared to axis sliders; details are provided in Appendix A.
@@ 5.2 Overlap Aware Modeling
Our differentiable framework automates this process by updating the axis weights to resolve overlaps efficiently. (See Fig. 10). Collision detection is performed by densely sampling points along all glyph outlines and any background objects. For each colliding point, we find the closest point on the other surface and its normal, defining a unilateral collision plane. Let $\mathbf{t}$ contain the curve indices and Bézier parameters for all points that need to be projected, and let $\mathbf{p}(\Theta,\mathbf{t})$ give the corresponding stacked positions under the current axis weights. Similarly, let $\mathbf{b}$ and $\mathbf{N}$ be the stacked closest points and normals. We define the collision energy as a differentiable
@@ 5.3 Physics Driven Kinetic Typography
For time integration, we adopt an iterative constraint projection approach inspired by projective dynamics [BML*14]. Unlike standard projective dynamics, however, we cannot rely on a global linear solve since the collision constraint projection is nonlinear. Instead, we iteratively minimize the combined elastic, kinetic, and collision energies using the Levenberg–Marquardt optimizer at each timestep.
@@ 5.4 Font Matching
using gradient descent. In practice, we use differentiable vector graphics rasterization [LLMR20] and employ an Adam optimizer (with 100 iterations), which reliably converges to visually accurate matches without additional tuning (Fig. 12). Since glyph geometry often contains overlapping elements, correct rasterization is obtained by specifying the nonzero fill rule in the differentiable SVG rasterizer, without requiring any special handling during optimization.
@@
5.5 Performance
We report representative runtime statistics for our applications to provide a qualitative sense of performance. All experiments were run using unoptimized PyTorch code on a MacBook Pro (M1 Pro, CPU only).
The average time per update call for interactive direct manipulation, with overlap-aware modeling enabled, is 0.024 s for Fig. Differentiable Variable Fonts(F), 0.049 s for Fig. Differentiable Variable Fonts(NT), and 0.033 s for Fig. 10 (“ello”).
Physics-driven kinetic typography is evaluated as an offline simulation. Fig. Differentiable Variable Fonts(O) requires 5.27 s for 1000 timesteps, while Fig. 11(stack) takes 9.09 s for 350 timesteps.
Font matching is an offline optimization task dominated by differentiable rasterization. Per-iteration runtimes are 0.37 s at resolution 256×256 (Fig. 12(A)), 1.77 s at 576×512 (Fig. 12(V)), and 2.21 s at 640×640 (Fig. Differentiable Variable Fonts(S)).
@@ 7 Conclusion
8 Acknowledgements
We thank Skef Iterum for feedback on our variable font formulation, Mengfei Liu for narrating the accompanying video, Zhecheng Wang, Chang Yue, Karran Pandey, Chenxi Liu, and Abhishek Madan for proofreading. Our research is funded in part by NSERC Discovery (RGPIN–2022–04680), the Ontario Early Research Award program, the Canada Research Chairs Program, a Sloan Research Fellowship, the DSI Catalyst Grant program and gifts by Adobe Inc.
We thank the designers and foundries of the variable fonts used in our figures: Recursive (Fig. 1; Fig. 12(c)), Angst (Fig. 2–3), Noto Sans Simplified Chinese (Fig. 3), Afacad Flux (Fig. 4–8; Fig. 11), Graduate (Fig. 9; Fig. 10(a)), Amstelvar (Fig. 10(b); Fig. 13), Mash (Fig. 10(c)), Shantell Sans (Fig. 11), Kamino (Fig. 12(a)), DynaPuff (Fig. 12(b)), and Movement (Fig. 13).
Appendix A User Study
To assess whether direct manipulation could be useful for manipulating variable fonts, we conducted a pilot user study with six novice users (5 male, 1 female), all of whom were using variable fonts for the first time. Participants were asked to manipulate a glyph initialized with a variable font’s default parameters to match a target glyph shape sampled from the font’s design space. These tasks were performed using both variable font axis sliders and our direct manipulation interface, which allows users to select, drag, and fix points on the glyph outline.
As a preliminary usability assessment, participants completed the System Usability Scale (SUS) for both interfaces. The direct manipulation interface obtained a mean SUS score of 74.17, compared to 55.41 for the slider-based interface. In an additional comparison question, five out of six participants reported that the direct manipulation interface better captured their intent overall. Qualitative feedback suggested that direct manipulation was perceived as more intuitive and easier to learn, particularly for complex glyphs or fonts with many axes, while sliders were preferred for precise refinement and backtracking.
We emphasize that this was a small-scale pilot study intended to provide preliminary insights rather than statistically conclusive evidence. As discussed in [HBF23] many human studies in computer graphics are performative rather than informative; ours is not an exception. Nevertheless, the results provide encouraging evidence that direct manipulation may offer a more intuitive interaction paradigm for variable font editing.